You will be using three different NaOH concentrations (0.1 M, 0.01M, and 0.001M) and a 4.0 x 10-6 M crystal violet stock solution. In all three cases one of the reactants is limiting and the other is in excess. Which reactant is limiting?

1)Cystal Violet

2) NaOH
Answer #1

Answer: If we look at the three concentrations of NaOH solution used, is 0.1 M = 1X10-1M; 0.01M= 1X10-2M and 0.001M= 1X10-3 M and the concentration of crystal violet stock solution is 4 X10-6 M = 0.000004 M which is very less as compared to NaOH concentration, therefore, crystal violet is the limiting reactant.